Alicia opens her piggy bank to find only nickles and dimes inside. She counts the coins and finds that there are 38 coins totaling $3.20. How many nickles were in the piggy bank, and how many dimes where in the piggy bank?
Found 3 solutions by stanbon, solver91311, richwmiller:
Answer by stanbon(75887)

Alicia opens her piggy bank to find only nickles and dimes inside.
She counts the coins and finds that there are 38 coins totaling $3.20.
How many nickles were in the piggy bank, and how many dimes where in the piggy bank?
----------------
Equations:
Quantity: n + q = 38
Value::: 5n +25q = 320 cents
----------------------------------
multiply thru the Quantity equation by 4 to get:
5n + 5q = 190
-----
Subtract that from the Value Equation and solve for "q":
20q = 130
q = 6.5
==================
Note: That value of q should be a whole number.
Check the data in your problem.
==================================
Cheers,
Stan H.

Answer by solver91311(24713)



Let represent the number of nickels. Let represent the number of dimes. Each nickel is worth 5 cents, so the value of the nickels in cents is . Likewise, the value of the dimes is cents. Finally, the total amount of money is $3.20 which is the same as 320 cents. Putting it all together you have:





If you solve the first equation for in terms of you get:



giving you an expression involving that is equal to and that can be substituted in place of in the other equation.



Now solve for .









So there are 26 dimes and therefore there must be nickels.

Check: 26 dimes are worth $2.60, and 12 nickels are worth $0.60. $2.60 plus $0.60 is $3.20. Checks.

Super Double Plus Extra Credit:

Solve the system using either the Elimination Method or Cramer's Rule.

John


Answer by richwmiller(17219)

Stan read quarters when it should be dimes